what 2 digit number exceeds its reversal by 20%?

Answer by ankor@dixie-net.com(22740)

what 2 digit number exceeds its reversal by 20%?
:
Let 10x + y = the original number
then
10y + x = the reversed number
:
10x + y = 1.2(10y + x)
10x + y = 12y + 1.2x
10x - 1.2x = 12y - y
8.8x = 11y
divide both sides by 11
.8x = y
:
Only one single digit value for x will give equal an integer value for y:
x=5
y = .8(5)
y = 4
:
54 is our two digit number
:
:
Check this
54 = 1.2(45)